Two cars leave towns 900 km apart at the same time and travel toward each other. One car's rate is 20 km per hour faster than the other's. If they meet in 5 hours, what is the rate of the faster car? Do not do any rounding.
Answer by TimothyLamb(4379)

a = b + 20
---
the closing speed of the two cars:
a + b
b + 20 + b
2b + 20
---
s = d / t
d = s * t
---
when the two cars meet they have traveled a combined distance of 900 km:
900 = (2b + 20)5
900 = 10b + 100
10b = 800
---
b = 80 kph
a = 100 kph
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php